OpenStudy (anonymous):

how do you solve In(4x)=3 when you don't know what x is?

OpenStudy (anonymous):

you know that \[ln(x)=y\] means \[e^y=x\] so\[ ln(4x)=3\] means \[e^3=4x\] \[x=\frac{e^3}{4}\]

OpenStudy (anonymous):

but is that your answer. or do you have to solve e^3/4?

OpenStudy (anonymous):

that is the answer unless you want a decimal approximation. if so use a calculator to get \[\frac{e^3}{4}=5.021\]

OpenStudy (anonymous):

okay. thats what i got. :)